# AC-PKAN: Attention-Enhanced and Chebyshev Polynomial-Based Physics-Informed Kolmogorov–Arnold Networks

### v1

@@ Appendix B Mathematical Proofs
$[J_{\ell}]_{k,i}\;=\;\sum_{n=0}^{N}C_{\ell,k,i,n}\,T_{n}^{\prime}\bigl{(}\tanh(x_{\ell,i})\bigr{)}\;\gamma_{\ell,i}.$ (48)
@@
$[\widetilde{J}_{\ell}]_{k,i}\;=\;\sum_{n=0}^{N}C_{\ell,k,i,n}\,T_{n}^{\prime}\bigl{(}\tanh(x_{\ell,i})\bigr{)},$ (49)
@@
each $\sigma_{i}(J_{\ell})$ cannot exceed its un-normalized counterpart $\sigma_{i}(\widetilde{J}_{\ell})$. For a fixed threshold $\epsilon>0$, let
@@
$\text{rank}_{\epsilon}(J_{\ell})\;\leq\;\text{rank}_{\epsilon}(\widetilde{J}_{\ell}).$ (54)
@@
B.3 Proof of Theorem 3.3
Theorem B.4 (Exponential Decay in Infinite Depth).
@@
By differentiating (56) w.r.t. $\mathbf{x}_{l}$, each layer’s Jacobian $J_{l}\in\mathbb{R}^{n\times n}$ has entries
@@
By Oseledec’s Multiplicative Ergodic Theorem [47], these limits exist almost surely for products of i.i.d. random matrices. In our Cheby1KAN setting, the layers’ Jacobians $J_{l}$ approximate a family of random matrices (the Jacobian entries being determined by i.i.d. Gaussian coefficients $C_{l,k,i,m}$), making the product $J_{\text{total}}$ amenable to the same analysis as in classical random matrix theory.
@@
When each $J_{l}$ is sufficiently close (in distribution) to an $n\times n$ Ginibre matrix with i.i.d. Gaussian entries, the Lyapunov exponents $\{\lambda_{i}\}$ match those of Ginibre ensembles, given by [46]:
@@
Since $\lambda_{i}<\lambda_{1}$ for $i\geq 2$ (because $\psi$ is strictly increasing and $n-i+1<n$), we have
[…]
@@
B.4 Proof of Proposition 3.4
Theorem B.5.
@@
Therefore, for any order $k>n$, the $k$-th derivative of $A^{(l)}$ is identically zero. Consequently, the $k$-th derivative of the first part of (78) is zero.
@@
$\left.\quad-\theta_{U}\left[\omega_{U,1}\cos(h_{0}\theta_{U}+b_{U})-\omega_{U,2}\sin(h_{0}\theta_{U}+b_{U})\right]\right),$ (89)
and the derivatives of any finite order of this term are also non-zero. Additionally, the third component of (78), $\tfrac{\partial U}{\partial x}$, is non-zero.
@@
so that the last two terms of (78) can be written as
@@
Consequently, the $k$-th derivatives of the remaining parts of (78) are non-zero, and thus the $k$-th derivatives of (77) are non-zero. Therefore, for any positive integer $N$, the derivative $\tfrac{\partial^{N}y}{\partial x^{N}}$ exists and is non-zero, establishing that AC-PKAN can approximate PDEs of arbitrarily high order. ∎
Remark: The property of possessing non-zero derivatives of any finite order with respect to the input $x$ specifically addresses enhancements in KAN variants rather than in MLP-based models. The fitting capability of KAN models relies on polynomial functions with learnable parameters. To ensure non-zero derivatives in the output, the original B-spline KAN incorporates an additional nonlinear bias function $b(x)$. In contrast, other KAN variants, such as Cheby1KAN, rely solely on polynomial bases, which inevitably result in zero derivatives when the order of differentiation exceeds the polynomial degree. Therefore, Proposition 3.4 was introduced to provide a theoretical guarantee for AC-PKAN’s ability to solve any PDE, analogous to how the universal approximation theorem theoretically establishes the universal fitting capability of neural networks.
B.5 Proof of Proposition 3.5
Definition B.6.
@@
Lemma B.7 (the Rank-nullity theorem).
@@
We may now, by the Steinitz exchange lemma, extend $K$ with $n-k$ linearly independent vectors $w_{1},\ldots,w_{n-k}$ to form a full basis of $V$.
@@
This is a contradiction to $\mathcal{B}$ being a basis, unless all $\alpha_{j}$ are equal to zero. This shows that $T(\mathcal{S})$ is linearly independent, and more specifically that it is a basis for Im $\text{Im }T$.
@@
Theorem B.8.
Let $\mathcal{N}$ be an AC-PKAN model with $L$ layers ($L\geq 2$) and infinite width. Then, the Jacobian matrix $J_{\mathcal{N}}(\boldsymbol{x})=\left[\frac{\partial\mathcal{N}_{i}}{\partial x_{j}}\right]_{m\times d}$ is full rank in the input space $\mathbb{R}^{d}$.
@@
Following the recursive relationship in Theorem  B.5 , $\frac{\partial\alpha^{(L)}}{\partial x}$ can be decomposed into a nonlinear combination of parameters across layers. Specifically, for any layer $l$, the derivative term $\frac{\partial\alpha^{(l)}}{\partial x}$ is generated through recursive operations involving parameters $C_{k,i,n}^{(l)}$, $\omega_{U},\omega_{V}$, $\theta_{U},\theta_{V}$, etc.
@@
$\frac{\partial y_{k}}{\partial x}=\mathbf{w}_{k}^{\top}\frac{\partial\alpha^{(L)}}{\partial x},\quad\frac{\partial y_{k^{\prime}}}{\partial x}=\mathbf{w}_{k^{\prime}}^{\top}\frac{\partial\alpha^{(L)}}{\partial x},$ (105)
@@
$\forall c_{i}\in\mathbb{R},\quad\sum_{i=1}^{d}c_{i}\mathbf{v}_{i}=0\implies c_{i}=0,\ \forall i$ (106)
@@
$\langle\mathbf{w}_{k},\mathbf{w}_{k^{\prime}}\rangle=\mathbf{w}_{k}\mathbf{w}_{k^{\prime}}^{\top}=\delta_{kk^{\prime}}\|\mathbf{w}_{k}\|^{2},\quad\forall k\neq k^{\prime}$ (107)
@@
$\mathbf{w}_{k}^{\top}\left(\sum_{i=1}^{d}\mathbf{v}_{i}\mathbf{v}_{i}^{\top}\right)\mathbf{w}_{k^{\prime}}=0\quad\forall k\neq k^{\prime}$ (110)
@@
Inductive Step: Let $P_{\mathcal{H}_{L-1}}^{\perp}$ be the orthogonal projection onto $\mathcal{H}_{L-1}^{\perp}$. Through Equations 76 and 79, we decompose:
@@
$\text{For }k\neq k^{\prime},\forall a,b\in\mathbb{R},\quad a\frac{\partial y_{k}}{\partial x}+b\frac{\partial y_{k^{\prime}}}{\partial x}=0\Rightarrow a=b=0.$ (115)
Consider the Jacobian matrix $J_{\mathcal{N}}(\boldsymbol{x})$ as a linear mapping $J_{\mathcal{N}}(\boldsymbol{x}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$. According to the rank-nullity theorem, we have:
$dim(ker(J_{\mathcal{N}}(\boldsymbol{x})))+rank(J_{\mathcal{N}}(\boldsymbol{x}))=d$ (116)
Theorem  B.5 guarantees that $rank(J_{\mathcal{N}}(\boldsymbol{x}))=min(d,m)$. Thus, the dimension of the kernel space is :
$dim(ker(J_{\mathcal{N}}(\boldsymbol{x})))=d-min(d,m).$ (117)
@@
  • When $m\geq d$: the Jacobian matrix has full column rank $rank(J_{\mathcal{N}}(\boldsymbol{x}))=d$, resulting in $ker(J_{\mathcal{N}}(\boldsymbol{x}))=\boldsymbol{0}$. $J_{\mathcal{N}}(\boldsymbol{x})$ is injective.
  • When $m<d$:the Jacobian matrix has full row rank $rank(J_{\mathcal{N}}(\boldsymbol{x}))=m$, resulting in $ker(J_{\mathcal{N}}(\boldsymbol{x}))=\boldsymbol{d-m}$, which means there exist $\boldsymbol{d-m}$ linearly independent non-zero vectors such that $J_{\mathcal{N}}(\boldsymbol{x})\boldsymbol{v}=\boldsymbol{0}$.
Let us exclude non-zero null vectors by contradiction. Assume there exists a non-zero vector $\boldsymbol{v}\neq\boldsymbol{0}\in\mathbb{R}^{d}$ such that . For any output component $\mathcal{N}_{i}$, we have
@@
According to Theorem  B.5 and Equation  115, the only solution is $\boldsymbol{v}=\boldsymbol{0}$, which contradicts the assumption. Therefore, the null space contains only the zero vector, i.e., $dim(ker(J_{\mathcal{N}}(\boldsymbol{x})))=\boldsymbol{0}$.
@@
$\text{rank}\left(J_{\mathcal{N}}(\boldsymbol{x})\right)<\min(d,m)\quad\forall\boldsymbol{x}\in\mathcal{M}.$ (119)
This implies that the image of the mapping $\mathcal{N}(\boldsymbol{x})$ is constrained to a lower-dimensional submanifold $\mathcal{S}\subset\mathbb{R}^{m}$, where:
$\dim(\mathcal{S})\leq\text{rank}\left(J_{\mathcal{N}}(\boldsymbol{x})\right)<\min(d,m).$ (120)
[…]

### v2

@@ A.2 Proof of Theorem 2
$J_{\ell}]_{k,i}\;=\;\sum_{n=0}^{N}C_{\ell,k,i,n}\,T_{n}^{\prime}\bigl(\tanh(x_{\ell,i})\bigr)\;\gamma_{\ell,i}.$ (49)
@@
$\widetilde{J}_{\ell}]_{k,i}\;=\;\sum_{n=0}^{N}C_{\ell,k,i,n}\,T_{n}^{\prime}\bigl(\tanh(x_{\ell,i})\bigr),$ (50)
@@
[…]
@@
[…]
@@
A.3 Proof of Theorem 3
Theorem 6 (Exponential Decay in Infinite Depth).
@@ A.3 Proof of Theorem 3
By differentiating equation 57 w.r.t. $\mathbf{x}_{l}$, each layer’s Jacobian $J_{l}\in\mathbb{R}^{n\times n}$ has entries
@@
By Oseledec’s Multiplicative Ergodic Theorem (Oseledets, 1968), these limits exist almost surely for products of i.i.d. random matrices. In our Cheby1KAN setting, the layers’ Jacobians $J_{l}$ approximate a family of random matrices (the Jacobian entries being determined by i.i.d. Gaussian coefficients $C_{l,k,i,m}$), making the product $J_{\text{total}}$ amenable to the same analysis as in classical random matrix theory.
@@
When each $J_{l}$ is sufficiently close (in distribution) to an $n\times n$ Ginibre matrix with i.i.d. Gaussian entries, the Lyapunov exponents $\{\lambda_{i}\}$ match those of Ginibre ensembles, given by (Newman, 1986):
@@
Since $\lambda_{i}<\lambda_{1}$ for $i\geq 2$ (because $\psi$ is strictly increasing and ni+1<$n$-i+1<n), we have
$e^{\lambda_{i}-\lambda_{1}}\;<\;1,\quad\forall\,i\geq 2.$ (66)
@@
Remark 1.
It is important to note that Theorem˜3 is proved under a random-coefficient model with i.i.d. Gaussian coefficients at initialization, and that trained networks do not necessarily satisfy independence or Gaussianity. For this reason, Theorem˜3 should be read as a motivational sufficient-condition result rather than a universal claim about learned models. Instead, for trained networks, the more relevant theoretical evidence is given by our deterministic layerwise rank bounds (Theorems 1 and 2), which are further supported by the empirical measurements reported in the paper.
A.4 Proof of Proposition 1
Theorem 7.
@@ A.4 Proof of Proposition 1
First, we clarify that the term "width" for a KAN denotes functional width, the number of Chebyshev basis functions per edge (N+1), not the number of neurons as used for MLPs.
@@
Therefore, for any order $k>n$, the $k$-th derivative of $A^{(l)}$ is identically zero. Consequently, the $k$-th derivative of the first part of equation 79 is zero.
@@
$\left.\qquad-\theta_{U}\left[\omega_{U,1}\cos(h_{0}\theta_{U}+b_{U})-\omega_{U,2}\sin(h_{0}\theta_{U}+b_{U})\right]\right),$ (90)
and the derivatives of any finite order of this term are also non-zero. Additionally, the third component of equation 79, $\tfrac{\partial U}{\partial x}$, is non-zero.
@@
so that the last two terms of equation 79 can be written as
@@
Consequently, the $k$-th derivatives of the remaining parts of equation 79 are non-zero, and thus the $k$-th derivatives of equation 78 are non-zero. Therefore, for any positive integer NN, the derivative $\tfrac{\partial^{N}y}{\partial x^{N}}$ exists and is non-zero, establishing that AC-PKAN can approximate PDEs of arbitrarily high order. ∎
Remark: The property of possessing non-zero derivatives of any finite order with respect to the input $x$ specifically addresses enhancements in KAN variants rather than in MLP-based models. The fitting capability of KAN models relies on polynomial functions with learnable parameters. To ensure non-zero derivatives in the output, the original B-spline KAN incorporates an additional nonlinear bias function $b(x)$. In contrast, other KAN variants, such as Cheby1KAN, rely solely on polynomial bases, which inevitably result in zero derivatives when the order of differentiation exceeds the polynomial degree. Therefore, Proposition 1 was introduced to provide a theoretical guarantee for AC-PKAN’s ability to solve any PDE, analogous to how the universal approximation theorem theoretically establishes the universal fitting capability of neural networks.
A.5 Proof of Proposition 2
Definition 1.
@@ A.5 Proof of Proposition 2
Lemma 2 (the Rank-nullity theorem).
@@
We may now, by the Steinitz exchange lemma, extend $K$ with nkn-k linearly independent vectors $w_{1},\ldots,w_{n-k}$ to form a full basis of $V$.
@@
This is a contradiction to $\mathcal{B}$ being a basis, unless all $\alpha_{j}$ are equal to zero. This shows that $T(\mathcal{S})$ is linearly independent, and more specifically that it is a basis for $\text{Im }T$.
@@
Theorem 8.
Let $\mathcal{N}$ be an AC-PKAN model with $L$ layers ($L\geq 2$) and infinite width. Then, the Jacobian matrix $J_{\mathcal{N}}(\bm{x})=\left[\frac{\partial\mathcal{N}_{i}}{\partial x_{j}}\right]_{m\times d}$ is full rank in the input space $\mathbb{R}^{d}$.
@@
Following the recursive relationship in Theorem 7 , $\frac{\partial\alpha^{(L)}}{\partial x}$ can be decomposed into a nonlinear combination of parameters across layers. Specifically, for any layer $l$, the derivative term $\frac{\partial\alpha^{(l)}}{\partial x}$ is generated through recursive operations involving parameters $C_{k,i,n}^{(l)}$, $\omega_{U},\omega_{V}$, $\theta_{U},\theta_{V}$, etc.
@@
$\frac{\partial y_{k}}{\partial x}=\mathbf{w}_{k}^{\top}\frac{\partial\alpha^{(L)}}{\partial x},\qquad\frac{\partial y_{k^{\prime}}}{\partial x}=\mathbf{w}_{k^{\prime}}^{\top}\frac{\partial\alpha^{(L)}}{\partial x},$ (106)
@@
$\forall c_{i}\in\mathbb{R},\qquad\sum_{i=1}^{d}c_{i}\mathbf{v}_{i}=0\implies c_{i}=0,\ \forall i$ (107)
@@
$\langle\mathbf{w}_{k},\mathbf{w}_{k^{\prime}}\rangle=\mathbf{w}_{k}\mathbf{w}_{k^{\prime}}^{\top}=\delta_{kk^{\prime}}\|\mathbf{w}_{k}\|^{2},\qquad\forall k\neq k^{\prime}$ (108)
@@
$\mathbf{w}_{k}^{\top}\left(\sum_{i=1}^{d}\mathbf{v}_{i}\mathbf{v}_{i}^{\top}\right)\mathbf{w}_{k^{\prime}}=0\qquad\forall k\neq k^{\prime}$ (111)
@@
Inductive Step: Let $P_{\mathcal{H}_{L-1}}^{\perp}$ be the orthogonal projection onto $\mathcal{H}_{L-1}^{\perp}$. Through Equations 77 and 80, we decompose:
@@
$\text{For }k\neq k^{\prime},\forall a,b\in\mathbb{R},\qquad a\frac{\partial y_{k}}{\partial x}+b\frac{\partial y_{k^{\prime}}}{\partial x}=0\Rightarrow a=b=0.$ (116)
Consider the Jacobian matrix $J_{\mathcal{N}}(\bm{x})$ as a linear mapping $J_{\mathcal{N}}(\bm{x}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$. According to the rank-nullity theorem, we have:
$dim(ker(J_{\mathcal{N}}(\bm{x})))+rank(J_{\mathcal{N}}(\bm{x}))=d$ (117)
Theorem 7 guarantees that $rank(J_{\mathcal{N}}(\bm{x}))=min(d,m)$. Thus, the dimension of the kernel space is :
$dim(ker(J_{\mathcal{N}}(\bm{x})))=d-min(d,m).$ (118)
@@
  • When $m\geq d$: the Jacobian matrix has full column rank $rank(J_{\mathcal{N}}(\bm{x}))=d$, resulting in $ker(J_{\mathcal{N}}(\bm{x}))=\bm{0}$. $J_{\mathcal{N}}(\bm{x})$ is injective.
  • When $m<d$:the Jacobian matrix has full row rank $rank(J_{\mathcal{N}}(\bm{x}))=m$, resulting in $ker(J_{\mathcal{N}}(\bm{x}))=\bm{d-m}$, which means there exist $\bm{d-m}$ linearly independent non-zero vectors such that $J_{\mathcal{N}}(\bm{x})\bm{v}=\bm{0}$.
Let us exclude non-zero null vectors by contradiction. Assume there exists a non-zero vector $\bm{v}\neq\bm{0}\in\mathbb{R}^{d}$ such that . For any output component $\mathcal{N}_{i}$, we have
@@
According to Theorem 7 and Equation 116, the only solution is $\bm{v}=\bm{0}$, which contradicts the assumption. Therefore, the null space contains only the zero vector, i.e., $dim(ker(J_{\mathcal{N}}(\bm{x})))=\bm{0}$.
@@
$\text{rank}\left(J_{\mathcal{N}}(\bm{x})\right)<\min(d,m)\qquad\forall\bm{x}\in\mathcal{M}.$ (120)
This implies that the image of the mapping $\mathcal{N}(\bm{x})$ is constrained to a lower-dimensional submanifold $\mathcal{S}\subset\mathbb{R}^{m}$, where:
$\dim(\mathcal{S})\leq\text{rank}\left(J_{\mathcal{N}}(\bm{x})\right)<\min(d,m).$ (121)
By Theorem 7, however, all first-order partial derivatives $\frac{\partial\mathcal{N}_{i}}{\partial x_{j}}\neq 0$. Specifically: (1)  $\forall\text{ direction }\mathbf{v}\in\mathbb{R}^{d}\setminus\{\bm{0}\}$ $\exists$ at least one output component $\mathcal{N}_{i}$ such that $\frac{\partial N_{i}}{\partial v}\neq 0$; (2) The infinite-width architecture of AC-PKAN ensures that the parameter space is dense in the $L^{2}$ function space. Consequently, the image set of the output mapping can densely cover any open set in $\mathbb{R}^{m}$.